The indentfirst package^{∗}

David Carlisle carlisle@cs.man.ac.uk

1995/11/23

Abstract

Make the first line of all sections etc., be indented by the usual paragraph indentation. This should work with all the standard document classes.

\if@afterindent L^{A}TEX uses the switch \if@afterindent to decide whether to indent after a section heading. We just need to make sure that this is always true.

1 ⟨∗package⟩

2 \let\@afterindentfalse\@afterindenttrue
3 \@afterindenttrue

4 ⟨/package⟩

Change History

v1.00

General: Initial version . . . . . . . . 1 v1.01

General: Re-issue for the new doc and docstrip . . . . . . . . . . . . . . 1 v1.02

General: Re-issue for LaTeX2e … 1 v1.03

General: Typo fixes in documentation . . . . . . . . . . . . . . . . . . . . 1

Index

Numbers written in italic refer to the page where the corresponding entry is described, the ones underlined to the code line of the definition, the rest to the code lines where the entry is used.

Symbols

\@afterindentfalse . . 2
\@afterindenttrue 2, 3

I

                         \if@afterindent . . . . \underline{1}

L \let . . . . . . . . . . . . . . 2

∗This file has version number v1.03, last revised 1995/11/23.